OpenStudy (anonymous):

Convert this to ft/s 600 inch/min

OpenStudy (anonymous):

\[\frac{600\text{ in}}{1\text{ min}}\times\frac{1\text{ ft}}{12\text{ in}}\times\frac{1\text{ min}}{60\text{ s}}=?\frac{\text{ft}}{\text{s}}\]

OpenStudy (anonymous):

600 in is 50 ft and 600 sec is 10 min